OpenStudy (anonymous):

Find the derivative of f(x) = negative 9 divided by x at x = -8.

OpenStudy (campbell_st):

write it in index form \[f(x) = -9x^{-1}\] now find the derivative lastly substitute x = -8 into the derivative

OpenStudy (mathstudent55):

\(\dfrac{d}{dx} x^n = n x^{n - 1} \)

OpenStudy (anonymous):

my answer choices are: 8/9 9/8 64/9 9/64

OpenStudy (anonymous):

when i did the equation you gave me i got 9/8 is that right??

OpenStudy (mathstudent55):

Can you first use the formula to find the derivative? Then we'll find the value.

OpenStudy (campbell_st):

well what does your derivative look like...?

OpenStudy (anonymous):

9/8 i did -9(-8)^-1

OpenStudy (campbell_st):

well remember you have to subtract 1 from the power of x so \[f'(x) = 9x^{-2}...or.....f'(x) = \frac{9}{x^2}\] now substitute x = -8

OpenStudy (anonymous):

9/64

OpenStudy (mathstudent55):

\(f(x) = -\dfrac{9}{x} \) \(f(x) = -9x^{-1} \) \(f'(x) = -1(-9)x^{-2} \) \(f'(x) = 9x^{-2} \) \(f'(x) = \dfrac{9}{x^2} \) \(f'(8) = \dfrac{9}{8^2} = \dfrac{9}{64} \) You are correct.

OpenStudy (anonymous):

thx

OpenStudy (mathstudent55):

wlcm